# DiffAdapt: Difficulty-Adaptive Reasoning for Token-Efficient LLM Inference

### v2

@@ B.4 Design Rationale
Appendix C Additional Experimental Details
C.1 Oracle Experiment Detailed Results
@@
C.2 Model-Specific Threshold Values
@@
C.3 Detailed Algorithmic Procedures
@@
C.4 Maximum Token Limits per Model and Benchmark
This subsection reports the maximum token limits used for each model on each benchmark and describes how they were determined. For each model–benchmark pair, we first ran the model under a generous cap (e.g., 32K tokens) to observe its longest response length in a less constrained setting. We then selected a nearby rounded integer as the per-benchmark max_tokens used in our analyses. This procedure standardizes evaluation across tasks and enables percentage-based truncation in Figures 5 and 6.
@@ C.4 Maximum Token Limits per Model and Benchmark
Appendix D Additional Overthinking Analysis Across Model Architectures
@@
D.1 DeepSeek-R1-Distill-Qwen-1.5B Overthinking Analysis
@@
D.2 Nemotron-Research-Reasoning-Qwen-1.5B Overthinking Analysis
@@
Appendix E Additional Oracle Analysis Results
@@
E.1 DeepSeek-R1-Qwen-7B Oracle Analysis
@@
E.2 Nemotron-1.5B Oracle Analysis
@@
E.3 DeepSeek-R1-Llama-8B Oracle Analysis
@@
E.4 Cross-Model Oracle Analysis Summary

### v3

@@ C.5 Hyperparameter Optimization
Appendix D Additional Experimental Details
D.1 Cross-Domain Generalization
To empirically demonstrate the robustness and transferability of DiffAdapt beyond pure math reasoning, we extended our evaluation to diverse out-of-domain benchmarks, including Minerva (scientific reasoning), GPQA (graduate-level domain knowledge), and MMLU-Pro (comprehensive general reasoning).
MMLU-Pro Results. We report the detailed zero-shot transfer results on MMLU-Pro in Table 9. By using the probe and thresholds trained solely on the DeepMath dataset, DiffAdapt consistently outperforms the fixed-strategy baseline by 3-7% across different token budgets and model architectures (DeepSeek-R1-Qwen/Llama). This confirms that the “difficulty signal” captured by our probe is generic and effectively transfers to unseen domains without re-training.
Table 9: Performance comparison on MMLU-Pro (OOD Generalization). DiffAdapt is applied zero-shot using probes trained on math data.
DeepSeek-R1-Qwen-7B DeepSeek-R1-Llama-8B
Token Budget DiffAdapt (%) Baseline (%) Improvement DiffAdapt (%) Baseline (%) Improvement
33.3% 32.02 28.21 +3.81% 22.14 17.50 +4.64%
50.0% 35.24 31.07 +4.17% 31.29 27.86 +3.43%
66.7% 35.00 31.07 +3.93% 33.74 31.79 +1.95%
83.3% 35.83 30.36 +5.47% 35.45 32.14 +3.31%
100% 36.48 30.71 +5.77% 39.90 32.50 +7.40%
D.2 Comparison with “When-to-Think” Baselines
We further compared DiffAdapt against specialized “when-to-think” methods like ThinkLess. Unlike these methods which typically require expensive two-stage training (SFT + RL), DiffAdapt is a training-free, plug-and-play approach for the LLM.
Setup. We applied DiffAdapt to the ThinkLess Stage-1 model (‘TL-1.5B-Warmup‘) and compared it against their fully trained Stage-2 RL model (‘TL-1.5B-RL‘).
Results. Table 10 presents the results on MATH500 and GSM8K.
  • Efficiency: On GSM8K, DiffAdapt consistently outperforms the RL baseline while using fewer tokens.
  • Cost-Effectiveness: On MATH500, while the RL model achieves higher peak accuracy, DiffAdapt outperforms the Warmup baseline by significant margins (+4-8%) and achieves competitive performance to the RL model in low-resource regimes using $\sim$35-50% fewer tokens, without requiring any RL training.
Table 10: Comparison against ThinkLess (TL) baselines. DiffAdapt is applied to the TL-1.5B-Warmup model.
MATH500 GSM8K
Budget TL-Warmup + DiffAdapt TL-1.5B-RL Analysis TL-Warmup + DiffAdapt TL-1.5B-RL Analysis
33.3% 64.4% (851 tok) 55.6% (1039 tok) +8.8% / Less Toks 67.0% (336 tok) 66.1% (409 tok) +0.9%
50.0% 67.6% (954 tok) 63.3% (1460 tok) +4.3% / -35% Toks 77.0% (357 tok) 72.0% (465 tok) +5.0%
72.2% 68.0% (973 tok) 69.2% (1775 tok) Comparable / -50% Toks 78.2% (361 tok) 74.3% (510 tok) +3.9%
100% 68.3% (1003 tok) 73.5% (2020 tok) RL peaks higher 78.6% (364 tok) 78.8% (573 tok) Comparable
D.3 Oracle Experiment Detailed Results
@@
D.4 Model-Specific Threshold Values
@@
D.5 Detailed Algorithmic Procedures
@@
D.6 Maximum Token Limits per Model and Benchmark
This subsection reports the maximum token limits used for each model on each benchmark and describes how they were determined. For each model–benchmark pair, we first ran the model under a generous cap (e.g., 32K tokens) to observe its longest response length in a less constrained setting. We then selected a nearby rounded integer as the per-benchmark max_tokens used in our analyses. This procedure standardizes evaluation across tasks and enables percentage-based truncation in Figures 3 and 4.
@@ D.6 Maximum Token Limits per Model and Benchmark
Appendix E Reasoning Integrity Analysis
A primary concern with efficiency-oriented reasoning methods is the potential risk of compromising reasoning integrity—specifically, whether aggressive token reduction leads to early truncation or logical gaps. To rigorously evaluate this, we conducted a blind, pairwise LLM-as-a-Judge study.
E.1 Experimental Setup
We randomly sampled $N=50$ queries from the GSM8K test set. For each query, we generated two responses:
  • System A (DiffAdapt): Our proposed method with adaptive strategy selection.
  • System B (Baseline): The standard Normal strategy (Temperature=0.8, full token budget).
We employed Qwen3-30B-A3B as an impartial judge. To ensure fairness, the evaluation was blind (model identities were anonymized) and pairwise (side-by-side comparison). The judge was explicitly instructed to evaluate based on logical completeness, coherence, and conciseness, and to penalize any instances of unjustified truncation.
E.2 Evaluation Prompt
The specific prompt used for the LLM-as-a-Judge evaluation is provided below. It explicitly asks the judge to focus on the preservation of coherent reasoning under token constraints.
LLM-as-a-Judge Prompt You will compare two systems on the same GSM8K math word problem.
Problem:
{problem}
Ground-truth solution (for verification only):
{ground-truth solution}
System A:
- Strategy: {strategy_A}
- Tokens used: {tokens_A}
- Final prediction: {Correct/Incorrect}
Reasoning trace:
{reasoning_trace_A}
System B:
- Strategy: {strategy_B}
- Tokens used: {tokens_B}
- Final prediction: {Correct/Incorrect}
Reasoning trace:
{reasoning_trace_B}
Decide which reasoning trace better preserves coherent, logically complete reasoning under tight token budgets. Explain your reasoning and output the winner (System A, System B, or Tie).
E.3 Results and Analysis
Table 14 summarizes the results of the blind evaluation.
Table 14: Blind pairwise comparison of reasoning quality (N=50) by Qwen3-30B Judge.
Outcome Count Percentage Judge’s Common Rationale
DiffAdapt Wins 38 76% “More direct,” “Avoids unnecessary repetition,” “Efficient logic”
Baseline Wins 6 12% “More detailed explanation” (in 5 cases), “Truncation” (in 1 case)
Tie 6 12% “Both reasoning paths are identical”
Frequency of Logical Failure We performed a manual failure analysis on the 6 cases where the Baseline won:
  • Subjective Preference (5 cases): The Baseline produced a more verbose explanation which the judge preferred, even though DiffAdapt’s response was correct and logically complete.
  • Truncation Error (1 case): Only a single instance (2%) involved actual logical failure due to aggressive token reduction (misclassified as Easy).
This low failure rate (2%) confirms that DiffAdapt’s fallback mechanisms (Normal strategy for ambiguous cases) effectively preserve reasoning integrity while significantly reducing computational cost.
Appendix F Additional Overthinking Analysis Across Model Architectures
@@
F.1 DeepSeek-R1-Distill-Qwen-1.5B Overthinking Analysis
@@
F.2 Nemotron-Research-Reasoning-Qwen-1.5B Overthinking Analysis
@@
Appendix G Additional Oracle Analysis Results
@@
G.1 DeepSeek-R1-Qwen-7B Oracle Analysis
@@
G.2 Nemotron-1.5B Oracle Analysis
@@
G.3 DeepSeek-R1-Llama-8B Oracle Analysis
@@
G.4 Cross-Model Oracle Analysis Summary